A rectangle with a perimeter of 30 centimeters is twice as long as it is wide. What is the area of the rectangle in square centimeters?
Found 2 solutions by stanbon, thejackal:
Answer by stanbon(75887)

A rectangle with a perimeter of 30 centimeters is twice as long as it is wide. What is the area of the rectangle in square centimeters?
--------
width = w
length = 2w
------
P = 2(length + width)
30 = 2(2w+w)
15 = 3w
width = 5 cm
length = 10 cm
-----
Area = 5*10 = 50 cm^2
-----------------------------
Cheers,
Stan H.
==================

Answer by thejackal(72)

perimeter P = 2w + 2l where w = width and l = length
we are told that the length is twice as long as it is wide which said differently means its 2 times the width
we can thus write the perimeter P as P = 2w + 2(2w)
so lets solve for w; 30 = 2(w) + 4(w); 30 = 6(w) hence w = 5 therefore l will 2*5 = 10
now to find the area with is w * l = 5 * 10 = 50 cm2